OpenStudy (moongazer):

If x grams of salt is mixed with y grams of sugar, then the percent of salt in the mixture is? please explain (meadal for good answer)

OpenStudy (anonymous):

Ok normally you would find the percentage in doing x/(x+y) * 100

OpenStudy (anonymous):

x = grams of salt y = grams of sugar x+y = total grams of mixture percent of salt = x / (x+y) * 100%

OpenStudy (anonymous):

thats exactly finding the substance you want over the total mass that you have Ex : you have 4 grams of sugar, 5 grams of salt you do 5/9 = 0.56 then multiply by 100 you get 56%